OpenStudy (anonymous):

A line passes through the point negative four comma one and has a slope of two-thirds. If you use this slope to generate another point on the line, what are the coordinates of that point?

OpenStudy (kira_yamato):

Point will be adding 1 to the x-coordinate of the original point and adding 2/3 to the y-coordinate of the original point. i.e. (-3 , 1 2/3)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@Kira_Yamato

OpenStudy (kira_yamato):

I think I don't get your question, but there are infinite number of points. y - 1 = 2/3 (x + 4) y - 1 = 2/3 x + 8/3 y = 2/3 x + 11/3

OpenStudy (anonymous):

it is telling you that on a graph there is a slope of two thirds what will be the next point after -4,1?

OpenStudy (anonymous):

@Kira_Yamato

OpenStudy (kira_yamato):

There can be many points, just plug in any value for x to get a corresponding value for y. It depends on your step size. Because the idea is for every 1 unit increase in x, there will be a 2/3 unit increase in y. So you can plug any value for x and you get the value for y and that's the point you're looking for :)

OpenStudy (anonymous):

but im so confused on this question can you PLEASE do it? :(

OpenStudy (anonymous):

@Kira_Yamato

OpenStudy (ranga):

You are given one point on the line: (-4, 1) Slope = 2/3 Use this slope to generate another point on the line Slope = 2/3 means if I move 3 units to the right and 2 units up the new point will be on the line. So the new point is: (-4 + 3, 1 + 2) = (-1, 3)

OpenStudy (ranga):

You can move multiples of 3 in the direction of x and the same multiple of 2 in the direction of y axis and they will all be on the line.

Directrix (directrix):

This is a diagram of what is going on.